### v1

@@ 5 Conclusion
We have introduced FLIP, a novel object-centric fovea-like input patching-based vision model that enhances scalable and data-efficient learning in complex visual scenes. FLIP implements a scale-invariant multi-resolution patch sampling mechanism, focusing high-resolution sampling around object centers using Gaussian distributions. As a result, FLIP effectively concentrates on critical regions while maintaining computational efficiency.
Our dynamic resolution mask prediction module and dedicated spatial attention mechanism, which separates perceptual and positional codes by applying positional embeddings exclusively to queries and keys, further improves segmentation accuracy. This approach aligns with principles well-known in cognitive modeling and enhances spatial selectivity without compromising perceptual content.
Experimental results on our newly introduced ObjaScale dataset reveal that current state-of-the-art segmentation models lack a critical feature: effectively segmenting very small objects. FLIP fills this gap by maintaining strong accuracy for tiny objects through its fovea-like, scale-invariant patch sampling. Across synthetic datasets and standard benchmarks such as Hypersim, KITTI-360, and OpenImages, FLIP achieves competitive Intersection over Union (IoU) scores with higher efficiency. These findings highlight FLIP’s potential for significant impact in computer vision applications that require both robust and highly data-efficient, generalizable, fine-grained object segmentations.

### v2

@@ 5 Conclusion
We have introduced FLIP, a novel vision model that fundamentally rethinks object segmentation through fovea-inspired input patching. By directly sampling multi-resolution patches centered on objects of interest, FLIP achieves state-of-the-art segmentation performance while using orders of magnitude fewer parameters than existing approaches.
Our key innovation lies in the scale-invariant patch sampling mechanism that adapts processing resolution to object characteristics, allocating computational resources where they matter most. This design enables FLIP-Large to achieve 80.33% mean IoU with only 96.6M parameters—outperforming SAM-H’s 75.41% IoU (641.1M parameters) while being 6$\times$ faster. Even our smallest model, FLIP-Tiny (0.51M parameters), surpasses all SAM variants with 78.24% mean IoU.
The introduction of ObjaScale reveals a critical limitation in current segmentation models: their inability to effectively handle very small objects in high-resolution scenes. FLIP addresses this gap through its fovea-like sampling, maintaining 87–89% IoU on ObjaScale where SAM variants achieve only 67–74%. This capability, combined with consistent performance improvements across Hypersim, KITTI-360, OpenImages, COCO, and LVIS, demonstrates FLIP’s potential for applications requiring precise segmentation of objects across extreme scale variations.
Looking forward, FLIP’s architectural efficiency and strong generalization open new possibilities for real-time object-centric vision systems. Its success suggests that biologically-inspired selective attention mechanisms can deliver both superior accuracy and computational efficiency, paving the way for more capable and sustainable computer vision applications in domains ranging from autonomous navigation to medical imaging.